# Chance Constrained PDE-Constrained Optimal Design Strategies Under High-Dimensional Uncertainty

### v1

@@ Acknowledgments
The authors gratefully acknowledge the financial support received from the U.S. National Science Foundation (NSF) CAREER Award CMMI-2143662. The authors express their appreciation for the helpful discussions with Dr. Peng Chen, from Georgia Institute of Technology, regarding the problem formulation and software libraries. Additionally, the authors would like to acknowledge the support provided by the Center for Computational Research at the University at Buffalo.
Appendix A Adaptive optimization
Input : Initial design parameter $d_{0}$, smoothing parameter $\omega_{0}$, penalty parameter $\gamma_{0}$, power parameters $\sigma_{\omega}$ and $\sigma_{\gamma}$, $\hat{f}=f_{\text{QUAD}}(d)$
Output : Design gradient $d_{opt}$
@@ Appendix A Adaptive optimization
2while $||d_{k}-d_{k-1}||\leq\mathcal{\epsilon_{\text{out}}}$ or $k<k_{\text{max}}$; $d_{k+1}=\text{Inexact Newton Conjugate-Gradient}(d_{k},\mathcal{J}(d),\nabla_{d}\mathcal{J}(d),\epsilon_{\text{in}})$.; Evaluate approximate chance $\hat{f}_{k+1}$ at $d_{k+1}$.; Update $\omega_{k+1}=\sigma_{\omega}\omega_{k}$, $\gamma_{k+1}=\sigma_{\gamma}\gamma_{k}$;
3end while
4Obtain the design gradient $d_{opt}$;
@@
The gradient-based scheme uses a continuation scheme that gradually increases the smoothing parameter $\omega$ and penalty parameter $\gamma$ to achieve convergence [25]. An initial value of the smoothing parameter $\omega_{0}$ and the penalty parameter $\gamma_{0}$ is defined for the continuation scheme. These parameters are updated at the end of each iteration using scaling parameters $\sigma_{\omega}$ and $\sigma_{\gamma}$ for smoothing and penalty parameters, respectively. For $k^{th}$ iteration , the updated smoothing parameter is $\omega_{k+1}=\sigma_{\omega}\omega_{k}$ and the updated penalty parameter is $\gamma_{k+1}=\sigma_{\gamma}\gamma_{k}$. In an outer loop, we update the parameters $\omega$ and $\gamma$, whereas an Inexact Newton Conjugate-Gradient (INCG) is applied in the inner loop to solve the optimization problem (51). The termination criterion for the outer loop is if the maximum number of iterations $k_{\text{max}}$ is reached. For each iteration, the approximated chance is computed through the quadratic approximation $f_{\text{QUAD}}$ of the chance function as given in (14). The output is the optimal value of design parameter $d_{\text{opt}}$. The algorithm for this adaptive optimization is given in Algorithm 1.

### v2

@@ Acknowledgments
The authors gratefully acknowledge the financial support received from the U.S. National Science Foundation (NSF) CAREER Award CMMI-2143662. The authors express their appreciation for Dr.Peng Chen, from Georgia Institute of Technology for valuable discussions on problem formulation and software libraries. Furthermore, the authors acknowledge the support provided by the Center for Computational Research at the University at Buffalo.
Appendix A Gradient and Hessian action of design objective and chance constraint
This section outlines the evaluations of the $m$-gradient, $m^{f}$-gradient, $m$-Hessian, and $m^{f}$-Hessian at the mean of the uncertain parameters $\bar{m}$, which are essential for constructing the quadratic Taylor approximations of the design objective and chance constraint, as described in Section 3.1.
$m$-gradient: The gradient $Q^{m}(\bar{m})$ acting in the direction $\tilde{m}=m-\bar{m}$ is expressed as [24],
$\langle\tilde{m},\bar{Q}^{m}\rangle=\langle\tilde{m},\partial_{m}\bar{Q}+\partial_{m}{\bar{r}^{Q}}\rangle,$ (31)
where $r^{Q}=r(\mathbf{u},\mathbf{v}^{Q},m,d)$ and the adjoint variable $\mathbf{v}^{Q}$ is considered as the Lagrange multiplier associated with the state equation with respect to the objective function $Q$. Here $(\bar{\cdot})$ is the quantities at the mean of $m$, and $(\tilde{\cdot})$ indicates directional variables referred to the variables that define the direction along which perturbation is considered to compute the change in the functional to obtain the derivative of the functional. In order to obtain the $m$-gradient, we need to obtain the multipliers $\mathbf{u}$ and $\mathbf{v}^{Q}$, which can be obtained using the state problem,
[…]
and the adjoint problem,
[…]
$m$-Hessian: The Hessian $Q^{mm}(\bar{m})$ acting on $\hat{m}$ can be expressed as [24],
$\begin{split}\langle\tilde{m},\bar{Q}^{mm}\,\hat{m}^{Q}\rangle=&\langle\tilde{m},\partial_{m\mathbf{v}}\bar{r}^{Q}\,\hat{\mathbf{v}}^{Q}+\partial_{m\mathbf{u}}\bar{r}^{Q}\,\hat{\mathbf{u}}^{Q}+\partial_{m\mathbf{u}}\bar{Q}\,\hat{\mathbf{u}}^{Q}+\partial_{mm}\bar{r}^{Q}\,\hat{m}^{Q}\\ +\partial_{mm}\bar{Q}\,\hat{m}^{Q}\rangle.\end{split}$ (34)
Thus, computing $m$-Hessian requires evaluation of the Lagrange multipliers $\hat{\mathbf{u}}^{Q}$ and $\hat{\mathbf{v}}^{Q}$ by solving the incremental state problem:
[…]
and the incremental adjoint problem:
Find $\hat{\mathbf{v}}^{Q}$, such that
$\langle\tilde{\mathbf{u}},\partial_{\mathbf{uv}}\bar{r}^{Q}\,\hat{\mathbf{v}}^{Q}\rangle=-\langle\tilde{\mathbf{u}},\partial_{\mathbf{uu}}\bar{r}^{Q}\,\hat{\mathbf{u}}+\partial_{\mathbf{uu}}\bar{Q}\,\hat{\mathbf{u}}^{Q}+\partial_{\mathbf{u}m}\bar{r}^{Q}\,\hat{m}^{Q}+\partial_{\mathbf{u}m}\bar{Q}\,\hat{m}^{Q}\rangle.$ (36)
$m^{f}$-gradient: The gradient of $f$ at $\bar{m}$ acting in the direction $\tilde{m}$ defined as [21]:
$\langle\tilde{m},\bar{f}^{m}\rangle=\langle\tilde{m},\partial_{m}\bar{f}+\partial_{m}\bar{r}^{f}\rangle$ (37)
The computation of $m^{f}$-gradient requires the lagrange multiplier $\mathbf{v}^{f}$ which is computed through the adjoint problem as:
Find $\mathbf{v}^{f}$, such that
$\langle\tilde{\mathbf{u}},\partial_{\mathbf{u}}\bar{r}^{f}\rangle=-\langle\tilde{\mathbf{u}},\partial_{\mathbf{u}}\bar{f}\rangle,$ (38)
$m^{f}$-Hessian: The action of Hessian of $f$ can be computed at $\tilde{m}$ as [21]:
$\begin{split}\langle\tilde{m},\bar{f}^{mm}\hat{m}^{f}\rangle=\langle\tilde{m},\partial_{m\mathbf{v}}\bar{r}^{f}\hat{\mathbf{v}}^{f}+\partial_{m\mathbf{u}}\bar{r}^{f}\hat{\mathbf{u}}^{f}+\partial_{m\mathbf{u}}\bar{f}\hat{\mathbf{u}}^{f}+\partial_{mm}\bar{r}^{f}\hat{m}^{f}+\partial_{mm}\bar{f}\>\hat{m}^{f}\rangle\quad.\end{split}$ (39)
Thus, the solution of $m^{f}$-Hessian requires the computation of Lagrange multipliers $\hat{\mathbf{u}}^{f}$ and $\hat{\mathbf{v}}^{f}$, by solving the incremental state problem,
[…]
and the incremental adjoint problem,
[…]
Also, in order to compute $m^{f}$-Hessian, we solve a generalized eigenvalue problem in (22) for a finite number of eigenvalues $N^{f}_{\mathrm{eig}}$.
Appendix B Gradient of design functional
The Lagrangian formalism is employed to derive the $d$-gradient by considering a meta-Lagrangian functional
$+\tau_{\gamma}(\mathbb{E}[l_{\omega}(f_{\text{QUAD}}(m,d))]-\alpha_{c})$
$+\sum^{N_{\mathrm{eig}}^{f}}_{j=1}\langle\left(\psi_{j}^{f}\right)^{*},(\bar{f}^{mm}-{\lambda_{j}^{f}}\mathcal{K}^{-1})\psi_{j}^{f}\rangle$
$+\sum^{N_{\mathrm{eig}}^{f}}_{j=1}\left(\lambda_{j}^{f}\right)^{*}(\langle{\psi_{j}}^{f},\mathcal{K}^{-1}{\psi_{j}}^{f}\rangle-1)+\sum^{N_{\mathrm{eig}}^{f}}_{j=1}\langle\left(\hat{\mathbf{v}}_{j}^{f}\right)^{*},\partial_{\mathbf{vu}}\bar{r}_{f}\,{\hat{\mathbf{u}}_{j}}^{f}+\partial_{\mathbf{v}m}\bar{r}_{f}\,{\psi_{j}}^{f}\rangle$
$+\sum^{N_{\mathrm{eig}}^{Q}}_{j=1}\langle\left(\psi_{j}^{Q}\right)^{*},(\bar{Q}^{mm}-{\lambda_{j}^{Q}}\mathcal{K}^{-1})\psi_{j}^{Q}\rangle$
$+\sum^{N_{\mathrm{eig}}^{Q}}_{j=1}\left(\lambda_{j}^{Q}\right)^{*}(\langle{\psi_{j}}^{Q},\mathcal{K}^{-1}{\psi_{j}}^{Q}\rangle-1)+\sum^{N_{\mathrm{eig}}^{Q}}_{j=1}\langle{\hat{\mathbf{v}}_{j}}^{{Q}^{*}},\partial_{\mathbf{vu}}\bar{r}_{Q}\,{\hat{\mathbf{u}}_{j}}^{Q}+\partial_{\mathbf{v}m}\bar{r}_{Q}\,{\psi_{j}}^{Q}\rangle$
where $(\cdot^{*})$ represents the adjoint variables, $(\hat{\cdot})$ denotes the incremental variables. The terms on the right hand side of (42), correspond to the following: the quadratic approximation of the design objective (18), the state and adjoint problems for chance constraint $f$ and design objective $Q$, the generalized eigenvalue problems for chance constraint $f$ and design objective $Q$, the incremental state and adjoint problems for chance constraint $f$ and design objective $Q$.
For obtaining the design gradient, we need to compute five sets of functional derivatives to solve for the corresponding ten sets of adjoint variables,
$\left(\mathbf{u}^{f}\right)^{*},\left(\mathbf{u}^{Q}\right)^{*},\left(\mathbf{v}^{f}\right)^{*},\left(\mathbf{v}^{Q}\right)^{*},\left(\psi_{j}^{f}\right)^{*},\left(\psi_{j}^{Q}\right)^{*},\left(\hat{\mathbf{u}^{f}_{j}}\right)^{*},\left(\hat{\mathbf{u}}^{Q}_{j}\right)^{*},\left(\hat{\mathbf{v}^{f}_{j}}\right)^{*},\left(\hat{\mathbf{v}^{Q}_{j}}\right)^{*}\}$ in (42), while the adjoint variable set for the eigenvalue set $\{{\lambda_{j}^{f}}^{*},{\lambda_{j}^{Q}}^{*}\}$ is not required due to its independency of the design parameter $d$.
Once we have defined the functional $\mathcal{L}_{\mathrm{QUAD}}$, we need to apply the condition that the first variation of Lagrangian as zero with respect to Lagrange multipliers for both design objective $Q$ and chance constraint function $f$. We start by equating the variation of (42) as zero with respect to $\{\lambda_{j}^{f}\}$,
$\{\left(\psi^{f}_{j}\right)^{*}\}=\frac{1}{2}\bigg{(}1+2\beta_{V}\,\lambda^{f}_{j}\bigg{)}\,\{\psi^{f}_{j}\},\quad j=1,\cdots,N^{f}_{\mathrm{eig}}.$ (43)
Next, we equate the first variation of (42) as zero with respect to $\mathbf{v}_{j}^{f}$ and $\mathbf{u}_{j}^{f}$ which gives the incremental state problem,
$\textit{Find $\hat{\mathbf{u}}^{f}$, such that}\quad\big{<}\tilde{\mathbf{v}},\partial_{\mathbf{vu}}\bar{r}^{f}\,\mathbf{u}^{f*}_{j}\big{>}=-\big{<}\tilde{\mathbf{v}},\partial_{\mathbf{v}m}\bar{r}^{f}\,\psi^{{f}^{*}}_{j}\big{>}.$ (44)
and linear adjoint problem:
Find $\left(\mathbf{v}^{f}\right)^{*}$, such that (45)
$-\langle\tilde{\mathbf{v}},\partial_{\mathbf{v}m}\bar{r}^{f}m^{f}\rangle.$
Similarly, we set the first variation of the (42) to zero for the Lagrange multipliers related to design objective $Q$. Firstly, we equate the first variation of (42) w.r.t eigenvalues $\{\lambda^{Q}_{j}\}$ to zero,
$\{{\psi^{Q}_{j}}^{*}\}=\frac{1}{2}\bigg{(}1+2\beta_{V}\,\lambda_{j}^{Q}\bigg{)}\,\{\psi_{j}^{Q}\},\quad j=1,\cdots,N_{\mathrm{eig}}^{Q}.$ (46)
Furthermore, we let the first variation of (42) as zero w.r.t $\mathbf{v}^{Q}_{j}$ and $\mathbf{u}^{Q}_{j}$ leading to the incremental state problem:
$\textit{Find $\hat{\mathbf{u}}^{Q}$, such that}\quad\big{<}\tilde{\mathbf{v}},\partial_{\mathbf{vu}}\bar{r}^{Q}\,\left(\mathbf{u}^{Q}_{j}\right)^{*}\big{>}=-\big{<}\tilde{\mathbf{v}},\partial_{\mathbf{v}m}\bar{r}^{Q}\,\left(\psi^{Q}_{j}\right)^{*}\big{>}.$ (47)
and the incremental adjoint problem:
$\begin{split}\textit{Find $\hat{\mathbf{v}}^{Q}$, such that}&\quad\\ \big{<}\tilde{\mathbf{u}},\partial_{\mathbf{uv}}\bar{r}^{Q}\,\left(\mathbf{v}^{Q}_{j}\right)^{*}\big{>}=-\big{<}\tilde{\mathbf{u}},\partial_{\mathbf{uu}}\bar{r}^{Q}\,{\left(\hat{\mathbf{u}}^{Q}_{j}\right)}^{*}+\partial_{\mathbf{uu}}\bar{Q}\,&{\left(\hat{\mathbf{u}}^{Q}_{j}\right)}^{*}+\partial_{\mathbf{u}m}\bar{r}^{Q}\,\left(\psi^{Q}_{j}\right)^{*}\big{>},\end{split}$ (48)
similar to (35) and (36), respectively. In combination with (46), we can express the adjoint variables for both the incremental state and adjoint problems separately for each eigenvector,
$\{\left(\hat{\mathbf{u}}\right)^{*}_{j}\}$ $=\frac{1}{2}\bigg{(}1+2\beta_{V}\,\lambda^{Q}_{j}\bigg{)}\,\{\hat{\mathbf{u}}^{Q}_{j}\},\quad j=1,\cdots,N_{\mathrm{eig}}^{Q},$ (49)
$\{\hat{\mathbf{v}}^{*}_{j}\}$ $=\frac{1}{2}\bigg{(}1+2\beta_{V}\,\lambda^{Q}_{j}\bigg{)}\,\{\hat{\mathbf{v}}^{Q}_{j}\},\quad j=1,\cdots,N_{\mathrm{eig}}^{Q}.$ (50)
Finally, we equate the first variation of the Lagrangian (42) to zero with respect to adjoint or state, we obtain the linear state problem:
Find ${\mathbf{u}^{Q}}^{*}$, such that
$\langle\tilde{\mathbf{v}},\partial_{\mathbf{vu}}\bar{r}^{Q}\,\left(\mathbf{u}^{Q}\right)^{*}\rangle=$ $-2\beta_{V}\langle\tilde{\mathbf{v}},\partial_{\mathbf{v}m}\bar{r}^{Q}\,(\mathcal{K}\partial_{m}\bar{r}^{Q})\rangle$ (51)
$-\sum^{N_{\mathrm{eig}}^{Q}}_{j=1}\langle\tilde{\mathbf{v}},\partial_{\mathbf{vuu}}\bar{r}^{Q}\,\hat{\mathbf{u}}^{Q}_{j}\,\left(\hat{\mathbf{u}}^{Q}_{j}\right)^{*}+\partial_{\mathbf{vu}m}\bar{r}^{Q}\,\psi^{Q}_{j}\,\left(\hat{\mathbf{u}}^{Q}_{j}\right)^{*}\rangle,$
and linear adjoint problem:
Find $\left(\mathbf{v}^{Q}\right)^{*}$, such that
$\langle\tilde{\mathbf{u}},\partial_{\mathbf{uv}}\bar{r}^{Q}\,\left(\mathbf{v}^{Q}\right)^{*}\rangle=$ $-\langle\tilde{\mathbf{u}},\partial_{\mathbf{u}}\bar{Q}\rangle-2\beta_{V}\langle\tilde{\mathbf{u}},\partial_{\mathbf{u}m}\bar{r}^{Q}\,(\mathcal{K}\partial_{m}\bar{r}^{Q})\rangle$ (52)
$+\partial_{\mathbf{uuu}}\bar{Q}\,\hat{\mathbf{u}}^{Q}_{j}\,\left(\hat{\mathbf{u}}^{Q}_{j}\right)^{*}+\partial_{\mathbf{uu}m}\bar{r}^{Q}\,\psi^{Q}_{j}\,\left(\hat{\mathbf{u}}^{Q}_{j}\right)^{*}\rangle.$
The computation of $d$-gradient means that we have to compute how the cost functional $\mathcal{J}_{\mathrm{QUAD}}$ changes with small variations in $d$. This can be done by computing the Fréchet derivative of $\mathcal{J}_{\mathrm{QUAD}}$ w.r.t $d$, which can be expressed as $D_{d}\mathcal{J}_{\mathrm{QUAD}}(d)$. However, the computation of this Fréchet derivative $D_{d}\mathcal{J}_{\mathrm{QUAD}}(d)$ becomes computationally prohibitive when the design parameter $d$ is high-dimensional. Hence, we compute the gradient through the partial derivative of the lagrangian $\mathcal{L}_{\mathrm{QUAD}}$ with respect to design parameter $d$, which is $\partial_{d}\mathcal{L}_{\mathrm{QUAD}}$, as it only requires the solution of lagrange multipliers, which are independent of the dimension of design parameter $d$. Also, since the total derivative of $\mathcal{J}_{\mathrm{QUAD}}$ along a random direction $\tilde{d}$ is equal to the partial derivative of lagrangian $\mathcal{L}_{\mathrm{QUAD}}$ w.r.t $d$, as it depends on design parameter only. Therefore, the $d$-gradient can be derived as,
@@ Appendix B Gradient of design functional
Input : Design objective (7) and weak form of the PDE (4).
Output : Gradient of the quadratic approximation of the design objective $\mathcal{J}_{\mathrm{QUAD}}(d)$ with respect to the design parameter $d$.
2
3Compute the $m$-gradient (31);
4
5Compute the $m$-Hessian (34);
6
7Compute the $m_{f}$-gradient (37);
8
9Compute the $m_{f}$-Hessian (39);
10
11Obtain the generalized eigenpairs using the double-pass randomized algorithm;
12
13Solve for $\mathcal{J}_{\mathrm{QUAD}}$ by (30);
14
15Find the solution to the linear state problem (51);
16
17Find the solution to the linear adjoint problem (45);
18
19Solve for the $d$-gradient (53)
Algorithm 1 Analytical design gradient
Appendix C Adaptive optimization
The gradient-based scheme uses a continuation scheme that gradually increases the smoothing parameter $\omega$ and penalty parameter $\gamma$ to achieve convergence [21]. An initial value of the smoothing parameter $\omega_{0}$ and the penalty parameter $\gamma_{0}$ is defined for the continuation scheme. These parameters are updated at the end of each iteration using scaling parameters $\Lambda_{\omega}$ and $\Lambda_{\gamma}$ for smoothing and penalty parameters, respectively. For $k^{th}$ iteration , the updated smoothing parameter is $\omega_{k+1}=\Lambda_{\omega}\omega_{k}$ and the updated penalty parameter is $\gamma_{k+1}=\Lambda_{\gamma}\gamma_{k}$. In an outer loop, we update the parameters $\omega$ and $\gamma$, whereas an Inexact Newton Conjugate-Gradient (INCG) is applied in the inner loop to solve the optimization problem (28). The termination criterion for the outer loop is if the maximum number of iterations $k_{\text{M}}$ is reached. For each iteration, the approximated chance is computed through the quadratic approximation $f_{\text{QUAD}}$ of the chance function as given in (9). The output is the optimal value of design parameter $d_{\text{opm}}$.
Input : Initial design parameter $d_{0}$, smoothing parameter $\omega_{0}$, penalty parameter $\gamma_{0}$, scaling parameters $\Lambda_{\omega}$ and $\Lambda_{\gamma}$, $\hat{f}=f_{\text{QUAD}}(d)$
Output : Optimal Design $d_{opm}$
1
2while $||d_{k}-d_{k-1}||\geq\mathcal{\epsilon_{\text{out}}}$ or $k<k_{\text{M}}$; $d_{k+1}=\text{Inexact Newton Conjugate-Gradient}(d_{k},\mathcal{J}(d),\nabla_{d}\mathcal{J}(d),\epsilon_{\text{in}})$.; Evaluate approximate chance $\hat{f}_{k+1}$ at $d_{k+1}$.; Update $\omega_{k+1}=\Lambda_{\omega}\omega_{k}$, $\gamma_{k+1}=\Lambda_{\gamma}\gamma_{k}$;
3end
4Obtain the optimal design $d_{opm}$;